# A Survey on Large Language Model based Human-Agent Systems

### v1

@@ 2 LLM-Based Human-Agent Systems
We define LLM-based human-agent systems as interactive frameworks where humans actively provide additional information, feedback, or control during interaction with an LLM-powered agent to enhance system performance, reliability and safety Feng et al. (2024); Shao et al. (2024); Mehta et al. (2024). The core idea is synergy: combining unique human strengths—like intuition, creativity, expertise, ethical judgment, and adaptability—with LLM agent capabilities such as vast knowledge recall, computational speed, and sophisticated language processing. LLM-HAS builds upon core LLM agent components but places critical emphasis on the human’s interactive prowess:
@@
    Provide Information / Clarification: Humans provide essential context, domain expertise, preferences, or resolve ambiguities, helping agents interpret situations more accurately Naik et al. (2025); Kim et al. (2025).
@@
    Provide Feedback / Error Correction: Humans evaluate agent outputs and provide feedback, ranging from simple ratings to complex critiques, demonstrations or corrections, effectively guiding agents’ adjustment Gao et al. (2024b); Dutta et al. (2024); Li et al. (2024b).
@@
    Take Control / Action: In high-stakes or sensitive scenarios (e.g., healthcare, privacy, or ethics), humans retain the authority to override, redirect, or halt agent actions, ensuring accountability, safety, and alignment with human values Chen et al. (2025); Natarajan et al. (2025); Xiao and Wang (2023).
Figure 1 provides a generalized overview of LLM-based human-agent systems. These systems operate within a defined Environment (e.g., physical world, simulation) that provides context and stimuli. Human & Agent Profiling characterize the participants’ roles and goals, and the agent’s core LLM engine augmented with capabilities like planning, memory, and tool use. Human Feedback can occur during different phases in various types and granularities. Human-Agent Interaction Types may be collaborative (most common), competitive, cooperative, or mixed. The Orchestration layer governs high-level coordination—choosing a task strategy (e.g., sequential one-by-one versus parallel simultaneous execution) and a temporal synchronization mode (real-time synchronous exchanges versus delayed asynchronous workflows) so that each actor acts at the right moment. The Communication layer specifies how information flows—defining message structure (centralized, decentralized, hierarchical) and mode (conversation, observation signals, or shared message pools). The effective interplay and configuration of these components, particularly various human feedback, are critical for tailoring the system to specific tasks and optimizing the overall system’s performance, reliability and safety.
@@ 3 Core Components
In this section, we dissect LLM-HAS Systems, discussing the five key aspects: environment & profiling, human feedback, interaction type, orchestration paradigm, and communication.
@@ 3.1 Environment and Profiling
Human & Agent Profiling. Human participants can be broadly categorized as lazy or informative users. Lazy users provide minimal guidance, typically offering evaluative feedback such as binary correctness or scalar rating. In contrast, informative users engage deeply by offering demonstrations, detailed guidance, refinements, or even taking over parts of the task Wang et al. (2024c); Han et al. (2025). On the other side, agents are profiled by their roles and capabilities—ranging from general assistants to specialized personas like mathematicians, engineers, doctors, or cleaning robots—each tailored to the specific demands of their operational context Guo et al. (2024a); Samuel et al. (2024).
@@ 3.2 Human Feedback
Type Evaluative User provides an assessment of the agent’s output quality, typically as binary assessment, scalar rating, or preference ranking. ① Easy to collect, scalable. ② Less specific signal for improvement. CoELA (Zhang et al., 2024a), MINT (Wang et al., 2024c), MetaGPT (Hong et al., 2023)
Corrective User offers edits or fixes to the agent’s behavior. ① Highly informative, clear signal for improvement. ② Higher user effort, often fine-grained & interactive. PARTNR (Chang et al., 2024), MindAgent (Gong et al., 2023), AI Chains (Wu et al., 2022)
@@
Implicit Inferred by the agent observing user actions or control signals, rather than explicitly stated or direct output modifications. ① Natural, unobtrusive collection. ② Ambiguous, requires careful interpretation. ReHAC(Feng et al., 2024), Attentive Supp. (Tanneberg et al., 2024a), A2C (Tariq et al., 2025)
@@
Granularity Coarse-grained / Holistic Single assessment/signal for an entire agent output, trajectory, or task outcome. ①  Simple for user, good for overall assessment ②  Obscures specific errors, less precise learning signal. AssistantX (Sun et al., 2024a), Help Feedback (Mehta et al., 2024), HRC DMP (Liu et al., 2024)
@@
Fine-grained / Segment-Level Feedback targeting specific parts of agent output, actions, or process. ①  Precise learning signal, crucial for debugging complex skills ②  Potentially higher user effort/burden. Collaborative Gym (Shao et al., 2024), MTOM (Zhang et al., 2024b), FineArena (Xu et al., 2025)
@@
Phase Initial Setup & Goal Definition Feedback provided task execution, configuring the agent system and defining the task, goals, constraints, and preferenc. ①  Proactive alignment, prevents costly errors, sets constraints ②  Requires upfront user input. AgentCoord (Pan et al., 2024a), GDfC (Wang et al., 2025c), HA Comm. (Bansal et al., 2024)
@@
During Task Execution Online, interactive feedback while the agent is actively performing the task, enabling real-time adaptation. ①  Enables real-time adaptation, crucial for dynamic/collaborative tasks ②  Requires responsive interfaces. InteractGen (Sun et al., 2024b), CowPilot (Huq et al., 2025), EasyLAN (Pan et al., 2024c)
@@
Post-Task Eval. & Refinement Feedback provided after task completion to assess outcomes and provide suggestions for immediate revision or future improvement. ①  Non-disruptive, good for aggregate data/offline learning ①  No impact on completed task. Auto Agent (Pan et al., 2024b), WebCanvas (Pan et al., 2024d), MineWorld (Guo et al., 2025)
@@
Table 1: Dimensions of Human Feedback in LLM-based human–agent systems. These dimensions include feedback type, granularity, and phase. For each dimension, a summary, key characteristics, trade-offs, and example works are provided for comparison.
@@
Human Feedback Type. We categorize human feedback as evaluative, corrective, guidance, and implicit feedback. (1) Evaluative Feedback provides an assessment of the agent’s output quality, typically as preference ranking, scalar rating, or binary assessment. A prime example is preference ranking, where users compare agent outputs, forming the basis of Reinforcement Learning from Human Feedback (RLHF) Chaudhari et al. (2024). Alternatively, platforms like Uni-RLHF Yuan et al. (2024) support scalar ratings or binary assessments. (2) Corrective Feedback offers direct edits or fixes to the agent’s behavior. For instance, the PRELUDE Gao et al. (2024a) framework learns latent preferences from user edits made to agent-generated text. (3) Guidance Feedback means the human proactively provides instructions, critiques, or demonstrations to shape the agent’s behavior. Agents like InteractGen Sun et al. (2024b), AutoManual Chen et al. (2024) can be bootstrapped using initial demonstrations, while methods like Self-Refine Choudhury and Sodhi (2025) employ iterative critiques and refinements to improve outputs. (4) Implicit Feedback is inferred by the agent observing user actions or control signals, rather than explicitly stated or direct output modifications. For example, an agent might learn user priorities by observing how a user adjusts control sliders in a system like VeriPlan Lee et al. (2025), or infer preferences by analyzing user behaviors like clicks and purchases in frameworks such as AgentA/B Wang et al. (2025a). This contrasts with corrective feedback where the user directly edits the output; here, the agent interprets the user’s independent actions or control choices.
@@
Human Feedback Phase. Human feedback can be incorporated at different phases of the LLM-agent pipeline. (1) Initial Setup & Goal Definition occurs before task execution, configuring the agent system and defining goals, such as setting coordination strategies (AgentCoord Pan et al. (2024a)) or critiquing plans before execution (Ask-before-Plan Zhang et al. (2024c)). (2) During Task Execution involves online, interactive feedback while the agent is actively performing the task, enabling real-time adaptation. Examples include interactive instruction editing (InstructEdit Wang et al. (2023)), mid-task refinements (Mutual Theory of Mind Zhang et al. (2024b), Collaborative Gym Shao et al. (2024)), online interventions (HG-DAgger Kelly et al. (2019)), or interpreting concurrent user actions (REVECA Seo et al. (2025)). (3) Post-Task Evaluation & Refinement happens after task completion to assess outcomes and provide feedback for immediate revision or future improvement. Frameworks like WebCanvas Pan et al. (2024d) and Organized Teams Guo et al. (2024b) apply feedback loops after initial generation for benchmarking or offline learning, while AdaPlanner Sun et al. (2023) archives successful plans post-task as skills for future use. Integrating feedback during execution is increasingly important for dynamic tasks requiring adaptation.
@@ 3.3 Human-Agent Interaction Types
Interaction types define how individuals communicate, exchange information, and take actions with one another. In LLM-HAS, interactions tend to be more dynamic and complex compared to multi-agent systems (MAS). This complexity arises from the various roles and responsibilities assigned to both human agents and those based on LLMs, necessitating a finer-grained framework to describe their collaborative behaviors. The following categorization highlights the three key interaction types: Collaboration, Competition and Coopetition. Base on the collaboration pattern, the collaboration can be partitioned into four fine-grained subtype, which will introduced in Section 3.3.1.
@@ 3.3.1 Collaboration
Collaborations are by far the most common interaction and foundational interaction, which involve humans and LLM-based agents working together to achieve a common goal. This partnership combines human creativity and contextual understanding with LLM-based agents to address challenges and improve the efficiency and quality of results (Vats et al., 2024; Du et al., 2024). Depending on the type of collaboration considered, it can be categorized into four main fine-granted subtypes: (1) Delegation & Direct Command (Kiewiet and McCubbins, 1991), (2) Supervision (Loganbill et al., 1982) (3) Cooperation (Rand and Nowak, 2013), and (4) Coordination (Turvey, 1990).
Delegation & Direct Command. In this interaction modality, a controlling party, usually a human, assigns explicit tasks to the LLM-based agent by providing clear and direct instructions. The agent is expected to execute these directives autonomously, or on the behalf of human, ensuring that responsibilities are well-defined and actions align with the system’s overarching objectives. Unlike supervision, where strategies can be dynamically adjusted in response to new situations, delegation involves providing instructions upfront. This means the agent follows a predetermined set of tasks rather than adapting to the situation. For instance, the investor specifies their risk preference to the agent executing the investment strategy (Xu et al., 2025), driver utter the command to LLM-based agent (Cui et al., 2024), and humans issue explicit action directives for LLM-based agent execution (Seo et al., 2025).
Supervision. Supervision is the process by which one party, usually a human operator, oversees, monitors and guides the actions of an LLM based agent. This involves real time evaluation and intervention to ensure the agent’s output aligns with established goals and quality standards. Supervision also encompasses setting alert thresholds and providing corrective inputs when deviations occur. By maintaining a continuous feedback loop between the human and the agent, supervision helps calibrate agent behaviour, catch and mitigate errors before they propagate and build confidence in the system. It also enables agents to handle routine tasks with increasing independence. For instance, agents seek human confirmation at critical moments (Shao et al., 2024), agents notify human to check whether the action is align (Liu et al., 2023b), teleoperator monitor the LLM generated motion plans (Liu et al., 2023a).
Cooperation. Cooperation refers to the voluntary and joint efforts of multiple parties to achieve agreed-upon goals. Unlike coordination, which focuses on organizing and aligning tasks, cooperation combines the various efforts and outcomes of different individuals and LLM-based agents toward a common objective. It emphasizes collective commitment, mutual assistance, and the pooling of resources to attain a shared result, thereby fostering a collaborative problem-solving environment. For instance, the human robot coordination in household activities Chang et al. (2024), cooperative embodied language agent (CoELA) (Zhang et al., 2024a), human designers collaborate with the LLM-based agent (Sharma et al., 2024).
Coordination. Coordination is the organized process of aligning and synchronizing the actions of multiple human and LLM-based agents to achieve a shared objective. The key idea behind coordination is to avoid conflict and bias in both humans and LLM-based agents to reach the final goal. It involves clear communication, strategic planning, and the intentional division of tasks, ensuring that individual efforts are harmonized and effectively integrated to support common goals. For instance, human and agents work in the shared workplace to complete interdependent tasks (Zhang et al., 2024b), human and agent integration for the adaptive decision-making (Sun et al., 2024b), AgentCoord for coordination work between human and agent (Pan et al., 2024a).
@@ 3.3.2 Competition
Competition is a form of interaction where participants aim to achieve their own goals, which often conflict with the objectives of others. In the LLM-HAS, competition emerges when agents or humans seek to enhance their personal performance or obtain resources, even if it negatively impacts collective results. In addition, competition also necessitates effective balancing mechanisms, like performance regulation or conflict resolution strategies, to prevent unproductive behaviors and ensure that the overall goals of the system remain intact. For instance, simulating the human and LLM-based agent social behaviors in the SOTOPIA framework (Zhou et al., 2024).

### v2

@@ 2 LLM-Based Human-Agent Systems
We define LLM-based human-agent systems as interactive frameworks where humans actively provide additional information, feedback, or control during interaction with an LLM-powered agent to enhance system performance, reliability, and safety Feng et al. (2024); Shao et al. (2024); Mehta et al. (2024). The core idea is synergy: combining unique human strengths—like intuition, creativity, expertise, ethical judgment, and adaptability—with LLM agent capabilities such as vast knowledge recall, computational speed, and sophisticated language processing. LLM-HAS builds upon core LLM agent components but places critical emphasis on the human’s interactive role and capabilities:
@@
    Providing Information / Clarification: Humans provide additional information that agents lack or cannot reliably infer, such as login credentials, payment details, domain expertise, constraints, or resolve ambiguities, helping agents interpret situations more accurately Naik et al. (2025); Kim et al. (2025).
@@
    Providing Feedback / Error Correction: Humans evaluate agent outputs and provide feedback, ranging from simple ratings to complex critiques, demonstrations or corrections, effectively guiding agents’ adjustment Gao et al. (2024b); Dutta et al. (2024); Li et al. (2024b).
@@
    Taking Control / Action: In high-stakes or sensitive scenarios (e.g., healthcare, privacy, or ethics), humans retain the authority to override, redirect, or halt agent actions, ensuring accountability, safety, and alignment with human values Chen et al. (2025); Natarajan et al. (2025); Xiao and Wang (2023).
Figure 1 provides a generalized overview of LLM-based human-agent systems. These systems operate within a defined Environment (e.g., physical world, simulation) that provides context and stimuli. Human & Agent Profiling characterizes the participants’ roles and goals, and the agent’s core LLM engine augmented with capabilities like planning, memory, and tool use. Human Feedback can occur during different phases in various types and granularities. Human-Agent Interaction Types may be collaborative (most common), competitive, cooperative, or mixed. The Orchestration layer governs high-level coordination, choosing a task strategy (e.g., sequential one-by-one versus parallel simultaneous execution) and a temporal synchronization mode (real-time synchronous exchanges versus delayed asynchronous workflows) so that each actor acts at the right moment. The Communication layer specifies how information flows, defining message structure (centralized, decentralized, hierarchical) and mode (conversation, observation signals, or shared message pools). The effective interplay and configuration of these components, along with various human feedback, are critical for tailoring the system to specific tasks and optimizing the overall system’s performance.
@@ 3 Core Components
In this section, we examine LLM-HAS through five core aspects: environment & profiling, human feedback, interaction type, orchestration paradigm, and communication. These dimensions provide a unified standard for analyzing existing work and guiding the design of future systems.
@@ 3.1 Environment and Profiling
Human & Agent Profiling. Human participants can be broadly categorized as lazy or informative users. Lazy users provide minimal guidance, typically offering evaluative feedback such as binary correctness or scalar rating. In contrast, informative users engage deeply by offering demonstrations, detailed guidance, refinements, or even taking over parts of the task Wang et al. (2024b); Liu et al. (2024b); Han et al. (2025). On the other side, agents are profiled by their roles and capabilities, which range from versatile general assistants to specialized experts in mathematics, engineering, medicine, or robotic cleaning, each adapted to the particular demands of its operational context Guo et al. (2024a); Samuel et al. (2024).
@@ 3.2 Human Feedback
Type Evaluative User provides an assessment of the agent’s output quality, typically as binary assessment, scalar rating, or preference ranking. ① Easy to collect, scalable. ② Less specific signal for improvement. EmoAgent (Qiu et al., 2025), MINT (Wang et al., 2024b), SOTOPIA (Zhou et al., 2024)
Corrective User offers edits or fixes to the agent’s behavior. ① Highly informative, clear signal for improvement. ② Higher user effort, often fine-grained & interactive. SymbioticRAG (Sun et al., 2025), SWEET-RL (Zhou et al., 2025), AI Chains (Wu et al., 2022)
@@
Implicit Inferred by the agent observing user actions or control signals, rather than explicitly stated or direct output modifications. ① Natural, unobtrusive collection. ② Ambiguous, requires careful interpretation. MTOM(Zhang et al., 2024b), Attentive Supp. (Tanneberg et al., 2024a), MineWorld (Guo et al., 2025)
@@
Granularity Coarse-grained / Holistic Single assessment/signal for an entire agent output, trajectory, or task outcome. ①  Simple for user, good for overall assessment ②  Obscures specific errors, less precise learning signal. AssistantX (Sun et al., 2024a), Help Feedback (Mehta et al., 2024), AXIS (Lu et al., 2024)
@@
Fine-grained / Segment-Level Feedback targeting specific parts of agent output, actions, or process. ①  Precise learning signal, crucial for debugging complex skills ②  Potentially higher user effort/burden. Collaborative Gym (Shao et al., 2024), Prison Dilemm (Jiang et al., 2025), FineArena (Xu et al., 2025)
@@
Phase Initial Setup & Goal Definition Feedback provided before task execution, configuring the agent system and defining the task, goals, constraints, and preference. ①  Initial and proactive alignment, prevents costly errors, sets constraints ②  Requires upfront user input. AgentCoord (Pan et al., 2024a), GDfC (Wang et al., 2025c), SMALL (Wang et al., 2024c)
@@
During Task Execution Online, interactive feedback while the agent is actively performing the task, enabling real-time adaptation. ①  Enables real-time adaptation, crucial for dynamic/collaborative tasks ②  Requires timely notification and responsive interfaces. InteractGen (Sun et al., 2024b), CowPilot (Huq et al., 2025), EasyLAN (Pan et al., 2024b)
@@
Post-Task Eval. & Refinement Feedback provided after task completion to assess outcomes and provide suggestions for future use. ①  Non-disruptive, good for aggregate data/offline learning ②  No impact on completed task. HRT-ML (Liu et al., 2024b), M3HF (Wang et al., 2025d), MAIH (Wang et al., 2024c)
@@
Table 1: Dimensions of Human Feedback in LLM-based human–agent systems, including feedback type, granularity, and phase. For each dimension, a summary, key characteristics, and example works are provided for comparison. A detailed overview of human feedback types and their subtypes is provided in our appendix (Table B.1).
@@
Human Feedback Type. We categorize human feedback as evaluative, corrective, guidance, and implicit feedback. (1) Evaluative Feedback provides an assessment of the agent’s output quality, typically as preference ranking, scalar rating, or binary assessment. A prime example is preference ranking, where users compare agent outputs, forming the basis of Reinforcement Learning from Human Feedback (RLHF) Chaudhari et al. (2024). Alternatively, platforms like Uni-RLHF Yuan et al. (2024) support scalar ratings or binary assessments. (2) Corrective Feedback offers direct edits or fixes to the agent’s behavior. For instance, the PRELUDE Gao et al. (2024a) framework learns latent preferences from user edits made to agent-generated text. (3) Guidance Feedback means the human proactively provides instructions, critiques, or demonstrations to shape the agent’s behavior. Agents like InteractGen Sun et al. (2024b), AutoManual Chen et al. (2024a) can be bootstrapped using initial demonstrations, while methods like Self-Refine Choudhury and Sodhi (2025) employ iterative critiques and refinements to improve outputs. (4) Implicit Feedback is inferred by the agent observing user actions or control signals, rather than explicitly stated or direct output modifications. For example, an agent might learn user priorities by observing how a user adjusts control sliders in a system like VeriPlan Lee et al. (2025a), or infer preferences by analyzing user behaviors like clicks and purchases in frameworks such as AgentA/B Wang et al. (2025a). This contrasts with corrective feedback, where the user directly edits the output; here, the agent interprets the user’s independent actions or control choices.
@@
Human Feedback Phase. Human feedback can be incorporated at different phases of the LLM-agent pipeline Wang et al. (2025d). (1) Initial Setup & Goal Definition occurs before task execution, configuring the agent system and defining goals, such as setting coordination strategies (AgentCoord Pan et al. (2024a)) or critiquing plans before execution (Ask-before-Plan Zhang et al. (2024c)). (2) During Task Execution involves online, interactive feedback while the agent is actively performing the task, enabling real-time adaptation. Examples include interactive instruction editing (InstructEdit Wang et al. (2023)), mid-task refinements (Mutual Theory of Mind Zhang et al. (2024b), Collaborative Gym Shao et al. (2024)), or online interventions (HG-DAgger Kelly et al. (2019)). (3) Post-Task Evaluation & Refinement happens after task completion to assess outcomes and provide feedback for future use. Frameworks like MAIH (Wang et al., 2024c) and EmoAgent (Qiu et al., 2025) apply feedback loops after initial generation for benchmarking or offline learning, while AdaPlanner Sun et al. (2023) archives successful plans post-task as skills for future use. Table 1 summarizes different dimensions of human feedback, key characteristics, and example work.
@@ 3.3 Human-Agent Interaction Types
Interaction types define how individuals communicate, exchange information, and take actions with one another. In LLM-HAS, interactions tend to be more dynamic and complex compared to multi-agent systems. This complexity arises from the various roles and responsibilities assigned to both human agents and those based on LLMs, necessitating a finer-grained framework to describe their collaborative behaviors. The following categorization highlights the three key interaction types: Collaboration, Competition, and Coopetition.
@@ 3.3.1 Collaboration
Collaborations are by far the most common interaction and foundational interaction, which involve humans and LLM-based agents working together to achieve a common goal. This partnership combines human creativity and contextual understanding with LLM-based agents to address challenges and improve the efficiency and quality of results (Vats et al., 2024; Du et al., 2024; Sun et al., 2025). Depending on the type of collaboration considered, it can be categorized into four main fine-grained subtypes (Figure 3): (1) Delegation & Direct Command (Kiewiet and McCubbins, 1991), (2) Supervision (Loganbill et al., 1982), (3) Cooperation (Rand and Nowak, 2013), and (4) Coordination (Turvey, 1990).
@@
Delegation & Direct Command. In this interaction modality, a controlling party, usually a human, assigns explicit tasks to the LLM-based agent by providing clear and direct instructions. The agent is expected to execute these directives autonomously or on behalf of humans, ensuring that responsibilities are well-defined and actions align with the system’s overarching objectives. Unlike supervision, where strategies can be dynamically adjusted in response to new situations, delegation involves providing instructions upfront. This means the agent follows a predetermined set of tasks rather than adapting to the situation. For instance, an investor specifies their risk preference to the agent executing the investment strategy like FineArena (Xu et al., 2025), or a driver utters the command to LLM-based agent like Drive as you Speak (Cui et al., 2024).
Supervision. Supervision is the process by which one party, usually a human operator, oversees, monitors, and guides the actions of an LLM-based agent. This involves real time evaluation and intervention to ensure the agent’s output aligns with established goals and quality standards. Supervision also encompasses setting alert thresholds and providing corrective inputs when deviations occur. By maintaining a continuous feedback loop between the human and the agent, supervision helps calibrate agent behavior, catch and mitigate errors before they propagate and build confidence in the system. It also enables agents to handle routine tasks with increasing independence. For instance, agents notify humans to verify alignment (Liu et al., 2023b), and teleoperators monitor the LLM-generated motion plans (Liu et al., 2023a).
Cooperation. Cooperation refers to the voluntary and joint efforts of multiple parties to achieve agreed-upon goals. This collaboration type combines the various efforts and outcomes of different individuals and LLM-based agents toward a common objective. It emphasizes collective commitment, mutual assistance, and the pooling of resources to attain a shared result, thereby fostering a collaborative problem-solving environment. For instance, the human robot coordination in household activities Chang et al. (2024), the cooperative embodied language agent (CoELA) (Zhang et al., 2024a), human designers collaborating with an LLM-based agent (Sharma et al., 2024).
Coordination. Coordination is the organized process of aligning and synchronizing the actions of humans and LLM-based agents to achieve a shared objective. Unlike cooperation, the key idea behind coordination is to avoid conflict and bias in both humans and LLM-based agents to reach the final goal. It involves clear communication, strategic planning, and the intentional division of tasks, ensuring that individual efforts are harmonized and effectively integrated to support common goals. For example, humans and agents work in a shared workspace to complete interdependent tasks (Zhang et al., 2024b), human–agent integration supports adaptive decision-making (Sun et al., 2024b), and the collaborative framework facilitates coordination between humans and agents (Pan et al., 2024a).
@@ 3.3.2 Competition
Competition is a form of interaction where participants aim to achieve their own goals, which often conflict with the objectives of others. In the LLM-HAS, competition emerges when agents or humans seek to enhance their personal performance or obtain resources, even if it negatively impacts collective results. In addition, competition also necessitates effective balancing mechanisms, like performance regulation or conflict resolution strategies, to prevent unproductive behaviors and ensure that the overall goals of the system remain intact. For instance, the SOTOPIA framework simulates social behaviors between humans and LLM-based agents (Zhou et al., 2024).